4. The following relations describe monthly demand and supply for a computer support ser-


vice catering to small businesses.


Q D = 3,000 - 10P


Q S = -1,000 + 10P


where Q is the number of businesses that need services and P is the monthly fee, in dollars.


a. At what average monthly fee would demand equal zero?


b. At what average monthly fee would supply equal zero?


c. Plot the supply and demand curves.


d. What is the equilibrium price/output level?


e. Suppose demand increases and leads to a new demand curve:


Q D = 3,500 - 10P


What is the effect on supply? What are the new equilibrium P and Q?


f. Suppose new suppliers enter the market due to the increase in demand so the new supply


curve is Q = –500 + 10 P. What are the new equilibrium price and equilibrium quantity?


g. Show these changes on the graph.

1
Expert's answer
2022-03-08T17:03:31-0500

a)

"QD=3,000-10P"


"QD=0"


Therefore;


"3,000-10P=0"


"10P=3,000"


"P=300"

Hence, the average monthly fee would demand equal to zero is 300


b)

"QS=-1,000+10P"

"QS=0"


Therefore;


"-1,000=10P=0"


"10P=1,000"


"P=100"


Hence, the average monthly fee would supply equal to zero is 100


c)




d)

"QD=QS"


"3,000-10P=-1,000+10P"


"3,000+1,000=10P+10P"


"4,000=20P"


"P=200"

200 is the equilibrium price.


"QD=3,000-10(200)"


"=3,000-2,000"


"=1,000"

Hence equilibrium quantity is 1,000


e)

"QD=3,500-10P"


"3,500-10P=-1,000+10P"


"4,500=20P"


"P=225"


Quantity is given by;


"QD=3,500-10(225)"


"=3,500-2,250"


"=1,250"


Hence supply increases.


f)


"QS=-500+10P"


"-500+10P=3,000-10P"

"20P=3,500"


"P=175"


Hence quantity will be;


"=-500+10(175)"


"=-500+1,750"


"=1,250"